Vandalay Industries has $30 million of debt, $10 million of preferred stock and $60 million of common stock outstanding. The market cost of debt is 8%, the cost of preferred is 9% and the cost of common equity is 14%. The firm has a 35% corporate tax rate. What is Vandalay's WACC?
Answer #1

The weighted average cost of capital (WACC) is found using the following equation

WACC = we re + wp rp + wd rd \times ( 1 - tax rate)

we = weight of equity = $ 60 million \div ( $ 30 + $ 10 + $ 60 ) million = 0.6

re = cost of equity = 14%

wp = weight of preferred capital = $ 10 million \div $ 100 million = 0.10

rp = cost of preferred capital = 9 %

wd = weight of debt = $ 30 million \div $ 100 million = 0.30

rd = cost of debt = 8 %

WACC = 0.6 \times 0.14 + 0.10 \times 0.09 + 0.30 \times 0.08 \times ( 1 - 0.35)

The company's WACC = 10.86 %